# Delving Deeper Into Astromorphic Transformers

### v2

@@ IV Results
Our astromorphic transformer is trained and tested on two different tasks. These include sentiment classification on the IMDB dataset [58] and image classification on the CIFAR10 dataset [59]. Appendix C in the supplementary material lists all the associated hyper-parameters. To date, there has not been any language or vision model that incorporates transformer architecture in a manner that is bioplausible and derived from neuron-astrocyte interactions. We report an accuracy of $88.7\pm 0.2\%$ for the sentiment classification task on the IMDB dataset, utilizing the non-linearity and relative positional information in the model. We also achieve $97.0\pm 0.2\%$ accuracy using an astromorphic vision transformer on the CIFAR10 dataset.
@@ IV-A Hardware Specifications & Dataset Details
The astromorphic transformer under consideration has been executed through the utilization of a GPU-based platform running PyTorch. The computational procedures were executed on a hardware configuration comprising a $12$ core Intel(R) Xeon(R) Platinum $8167M$ CPU @ 2.00GHz, in conjunction with an NVIDIA® Tesla® V100 GPU possessing $16384$ Mega Bytes of graphics memory. While simulating the model on the IMDB dataset, the model uses 1.7 GB of memory whereas it uses 4.4 GB of memory on the CIFAR10 dataset.
The IMDB dataset, created by Stanford University, contains $50,000$ movie reviews (each 250-500 words long) labeled as positive or negative. The data is presented in a tab-delimited format, containing fields for movie title, review text, and sentiment label. The CIFAR-10 dataset is a collection of computer vision data, comprising $60,000$ color images with dimensions of $32\times 32$ pixels. These images are categorized into 10 distinct classes, with each class containing 6,000 images.
@@ IV-B Network Architecture
In the Sentiment classification task, a transformer with an encoder-based architecture is employed. We use a similar architecture as depicted in the Fig. S1 with 1 encoder layer and 4 headed self-attention. We use the combination of GloVe word embeddings with one-hot positional encoding as input tokens. A dropout layer is employed after the addition of residual connection just before the layer normalization.
@@
Meanwhile, for image classification, a convolutional and self-attention based neural network (CoAtNet [60]) was utilized to implement our proposed innovations. CoAtNet transformer architecture can best utilize the effectiveness of relative positional information when trained on vision task focused datasets. Furthermore, CoAtNet leverages the capabilities of Convolutional Neural Networks (CNNs) and transformers to enhance its performance. Our architecture implements a $3\times 3$ convolution before creating patches from the images. We use a patch size of $2$ on the image inputs and employ an 8-layered encoder stack with 8-headed self-attention. In addition to the original CoAtNet architecture, we add additional non-linearity and average pooling after encoder layers.
@@
For both the architectures, AdamW optimizer and Cross Entropy Loss are used to learn the weights of the neural network. For IMDB dataset tokenization, we used the Glove pretrained word vectors with 840B tokens, 2.2M vocab (cased), 300D vectors.
IV-C Performance Evaluation
Table I compares performance of our astromorphic transformer against neuromorphic as well as other non-bioplausible models on the IMDB and CIFAR10 datasets. Our model outperforms RNNs, LSTMs, and GRUs on the IMDB dataset and other bioplausible neuromorphic architectures. However, we did not compare with existing pretrained transformers (BERT, GPT, etc.) because these models are pretrained using non-bioplausible self-attention and have customized pretrained learnable embeddings for natural language processing, whereas we use fixed GloVe embeddings [61] that are simple enough to implement in our astromorphic attention-based transformer architecture. We also test the competitiveness of astromorphic attention in vision transformer architecture. As shown in Table I, our astromorphic vision transformer model competes with artificial and outperforms bioplausible models on the CIFAR10 dataset.
IV-D Ablation Studies
IV-D1 Learning speed
The impact of the two primary contributions of our study is illustrated in Fig. 4. It is evident from the Fig. that the absence of relative positional information leads to a decrease in the learning speed of the model. Measuring the number of epochs to reach $85\%$ ($85\%$) accuracy on the IMDB (CIFAR10) dataset, we find that: (i) with both of our contributions (non-linearity and relative positional information), it takes only $16\pm 2$ ($45\pm 3$) epochs, whereas (ii) with only the non-linearity included, it takes $22\pm 2$ ($111\pm 4$) epochs, and (iii) with only relative positional information, it takes $43\pm 4$ ($215\pm 2$) epochs, and (iv) without both, it takes $54\pm 4$ ($215\pm 2$) epochs to reach the specified accuracy.
IV-D2 Final accuracy
We implement the theoretical model proposed by [18] to see how it compares without the bioplausibilities proposed herein and find an accuracy of $88.0\pm 0.2\%$ and $87.4\pm 0.1\%$ on the IMDB and CIFAR10 datasets, respectively. This implementation does not incorporate the non-linearity and relative positional information that are integral to our model, which draws inspiration from the dynamics between neurons and astrocytes, thereby underscoring the strong need to consider such bioplausibilities in modeling astromorphic transformers. Our astromorphic transformer also competes with an iso-architecture standard non-bioplausible transformer with softmax attention (see “Transformer” entry in Table 1).
@@ V Conclusion
This study aims to provide a detailed explanation of the development of an astromorphic transformer using bioplausible self-attention. The inspiration for this approach is drawn from the dynamics of bidirectional communication between astrocytes and neurons in a tripartite synapse. In this study, we exploit the intrinsic Presynaptic and Hebbian plasticity of the astrocyte to incorporate two novel advancements into the transformer architecture. Subsequently, we demonstrate the effectiveness of our model on two distinct machine learning tasks. The performance of Astromorphic Transformers on sentiment and image classification tasks is presented for the first time on the IMDB and CIFAR10 datasets and is found to perform at par with an iso-architecture non-bioplausible transformer model. Implementation of Astromorphic Transformers on neuromorphic hardware has the potential of enabling scalable, robust, power and energy efficient computing paradigms that are applicable to a broad range of sequential processing tasks - a significant shift from the huge computational requirements of conventional deep learning solutions like Large Language Models.
Future research in this area could involve the exploration of astrocyte-to-astrocyte communication through gap junctions. This approach may enable the parallelization of Presynaptic and Hebbian plasticity, potentially addressing the limitation of a slower calcium response in comparison to neuron activation. In addition, it is possible to modify our model to accommodate large language models such as BERT and GPT. This can be achieved by incorporating the functionalization of relative positional information with the customized embedding of these models from a neuromorphic perspective. Further, quantifying the energy efficiency benefits of implementing Astromorphic Transformers on neuromorphic hardware can be an area of exploration.
@@ V-A Appendix A: Transformer Illustration
Transformer employs only encoder architecture for performing sentiment classification and vision tasks. As illustrated in Fig. S5, the input embeddings combined with the positional embeddings enter the encoder layer as input tokens. Then key, query and value are generated from the input tokens and after performing self attention on them, a skip connection is added from the inputs and the overall result is layer normalized and fed into a feed forward network. Another add & normalize operation is performed before feeding them into a linear layer and softmax function to obtain the output probabilities. Sometimes, there might be more than one encoder layer. For different tasks, like machine translation and text generation, encoder-decoder or only decoder type architecture of transformer can be used. However, independent of the type of architecture needed, self-attention is a fundamental mechanism defining transformer operation. The “Multi-Head Attention” Block has self-attention modules parallelly running on different heads obtained from the input embeddings.
@@ V-B Appendix B: Memory Efficient Linearized Attention
Linearized attention involves utilizing feature maps with lower complexity levels, thereby minimizing memory and computational resources. As discussed before, the introduction of non-linearity in biological neurons is facilitated by an activation function ($elu(X)$) in the hidden layer neurons. The utilization of the positive similarity feature map, denoted by $\phi(.)$, characterizes the implementation of the linearized attention operation as described in the work by (Katharopoulos et al. 2020). The utilization of linearized feature attention results in a reduction of computational complexity associated with the computation of the attention score. The advantage of utilizing feature map based linearized attention over the conventional distribution function (i.e. softmax) based self-attention is demonstrated in Fig. S6.
Usual self-attention employs softmax as the distribution function, however, the dot product of Query ($Q$) and Key ($K$) poses an $O(N^{2}D)$ complexity which overuses the memory. Whereas, in linearized attention the computation multiplies Key ($K$) and Value ($V$) yielding an $O(ND^{2})$ complexity, properly utilizing the memory and computational efficiency when token length ($N$) is larger than internal embedding size ($D$). Thus, the employment of the linearized feature map, denoted as $\phi(.)$, enables the realization of memory efficient linearized transformer.
TABLE S2: Hyper-parameter list for the IMDB dataset.
Hyper-parameters
Value
Tested Range
$\alpha$ (Non-linearity parameter)
$0.25$
$0.15$-$0.4$
Model size (D)
$128$
$32$-$256$
(Scaling parameter $m$)
$500$D
($1$-$1000$)D
Model size
$128$
$32$-$256$
Sequence length
$400$
$100$-$600$
Learning rate
$0.00001$
$41e-$ 4 - $61e-$ 6
Batch size
$64$
$8$-$256$
Number of epochs
$120$
$40$-$150$
Number of heads
$4$
$1$-$12$
TABLE S3: Hyper-parameter list for the CIFAR10 dataset.
Hyper-parameters
Value
Tested Range
$\alpha$ (Non-linearity parameter)
$0.25$
$0.15$-$0.4$
Model size (D)
$256$
$64$-$512$
(Scaling parameter $m$)
$500$D
($1$-$1000$)D
Patch size
$2$
$2$-$4$
Learning rate
$0.001$
$21e-$ 2 - $41e-$ 4
Batch size
$32$
$8$-$128$
Number of epochs
$250$
$80$-$300$
Number of heads
$8$
$2$-$12$
@@ V-C Appendix C: Hyper-parameter Details
The hyper-parameters for the encoder-type astromorphic transformer for the IMDB dataset are listed in Table S2 in the supplementary material. The hyper-parameters for the encoder-type convolution and astromorphic self-attention based transformer for the CIFAR10 dataset are listed in Table S3.

### v3

@@ IV Results
Our Astromorphic Transformer is trained and tested on three different tasks: sentiment classification on the IMDB dataset [65], image classification on the CIFAR10 dataset [66], and language modeling on the Wikitext-2 dataset [67]. To date, there has not been any language or vision model that incorporates transformer architecture in a bioplausible manner derived from neuron-astrocyte interactions. We report the following performance for each task:
  • $88.7\pm 0.2\%$ for sentiment classification on the IMDB dataset, utilizing the non-linearity and relative positional information in the model.
  • $97.0\pm 0.2\%$ accuracy for image classification using an astromorphic vision transformer on the CIFAR10 dataset.
  • A perplexity of $33.8$ for language modeling on the Wikitext-2 dataset using one decoder layer with 6-headed self-attention.
@@ IV-A Hardware Specifications & Dataset Details
The Astromorphic Transformer has been executed on a GPU-based platform running PyTorch. The hardware setup consists of a $12$-core Intel(R) Xeon(R) Platinum $8167M$ CPU @ 2.00GHz, with an NVIDIA® Tesla® V100 GPU (16 GB memory).
The IMDB dataset contains $50,000$ movie reviews labeled as positive or negative, with each review spanning between 250 and 500 words. The CIFAR10 dataset comprises $60,000$ color images categorized into 10 distinct classes, with each image having dimensions of $32\times 32$ pixels. The Wikitext-2 dataset is a widely used benchmark for language modeling and natural language generation tasks, containing over two million words with carefully preserved capitalization, punctuation, and rare terms.
@@ IV-B Network Architecture
For the sentiment classification task on the IMDB dataset, we employ a transformer with an encoder-based architecture, using one encoder layer with 4-headed self-attention. A dropout layer is applied after the residual connection before layer normalization.
@@
For the image classification task on the CIFAR10 dataset, we implement the CoAtNet transformer architecture, combining convolutional neural networks (CNNs) with self-attention. The architecture employs a $3\times 3$ convolution before creating patches, followed by an 8-layered encoder stack with 8-headed self-attention. A patch size of $2$ is used on the image inputs. In addition to the original CoAtNet architecture, we add non-linearity and average pooling after the encoder layers.
For the language modeling task on the Wikitext-2 dataset, we use a decoder-based transformer architecture with one decoder layer and 6-headed self-attention. The embeddings for this task are generated using the GPT-2 tokenizer, and they are passed through an astromorphic decoder block with layer normalization and dropout before the final output.
@@
For all tasks, we use AdamW optimizer and Cross Entropy Loss for training. For tokenization, the GloVe pretrained word embeddings [69] are used with 840B tokens, 2.2M vocab (cased), and 300D vectors for the IMDB dataset, and for CIFAR10 dataset, pixels are converted to sequences by using a patch size of 2, while GPT-2 tokenizers are used for the Wikitext-2 dataset. All hyperparameters for these architectures are detailed in Appendix C.
IV-C Performance Evaluation and Comparative Analysis
Table I compares the performance of our Astromorphic Transformer against both bioplausible neuromorphic models and non-bioplausible models on the IMDB, CIFAR10, and Wikitext-2 datasets. We chose not to compare our results against pretrained transformer models like BERT and GPT, as these architectures leverage pretraining on large datasets and use non-bioplausible self-attention mechanisms with highly customized learnable embeddings. In contrast, our model is trained from scratch, making it simpler to implement the astromorphic attention in the transformer framework while still achieving competitive results. The results, summarized in Table I, demonstrate the model’s effectiveness compared to both bioplausible and non-bioplausible architectures.
IV-C1 Sentiment Classification on IMDB
On the IMDB dataset, the Astromorphic Transformer achieves an accuracy of 88.7%, outperforming several traditional models, including RNNs, LSTMs, and SNNs. It also exceeds the accuracy of our implementation of Kozachkov et al.’s proposal [18]: Linearized Transformer inspired from astrocytes by 0.3%, which lacks the astrocytic non-linearity and relative positional encoding central to our model’s design.
IV-C2 Image Classification on CIFAR10
In the CIFAR10 dataset, the Astromorphic Transformer attains an accuracy of 97.0%, outperforming the Linearized Transformer inspired from astrocytes [18] which achieves 96.8%. Both models deliver high performance, but the Astromorphic Transformer shows a slight edge, benefiting from enhanced feature learning enabled by astrocytic non-linearity and positional encoding, especially in complex image classification tasks.
IV-C3 Language Modeling on Wikitext-2
For language modeling, the Astromorphic Transformer demonstrates superior performance with a perplexity of 33.8 on the Wikitext-2 dataset. In contrast, Linearized Transformer inspired from astrocytes [18] fails to converge due to gradient explosion, yielding a perplexity of 123.3. This stark difference highlights the necessity of incorporating astrocytic non-linearity for stabilizing training and improving model generalization in sequential tasks.
Figure 4: Comparative analysis to examine the effects of incorporating non-linearity and relative positional encoding in the Astromorphic Transformer on the (a) IMDB and (b) CIFAR10 datasets. The yellow and purple lines in Fig. (b) are nearly overlapping. The purple curve represents the result of the model formulated by [18] which performs subpar in both datasets compared to our proposed Astromorphic Transformer (blue curve).
IV-C4 Ablation Studies in Learning Speed
The impact of astrocytic non-linearity and relative positional encoding on learning speed is evident in both datasets. On the IMDB dataset, the Astromorphic Transformer reaches 85% accuracy in 16$\pm$2 epochs. Without positional encoding, this increases to 22$\pm$2 epochs, and without both features, it takes 54$\pm$4 epochs. These results reflect a 27.3% faster convergence with positional encoding and a 70.4% improvement when both features are included. In the CIFAR10 dataset, the model reaches 85% accuracy in 45$\pm$3 epochs. Without positional encoding, it requires 111$\pm$4 epochs, and without either feature, it stretches to 215$\pm$2 epochs, demonstrating a 59.5% faster convergence with positional encoding and a 79.1% speed improvement with both features. For Wikitext-2 dataset, the Astromorphic Transformer converges to the lowest perplexity within $40\pm 5$ epochs, whereas the Linearized Transformer inspired from astrocytes [18] fails to converge, and the model crashes due to gradient explosion after a few epochs. Fig. 4 makes it evident that our model significantly outperforms the Linearized Transformer inspired from astrocytes [18] in terms of both learning speed and final performance, demonstrating the crucial role of bioplausible features in improving the convergence and generalization of transformer models.
@@ V Conclusion
The Astromorphic Transformer advances the frontier of bioplausible computing by effectively leveraging neuron-astrocyte interactions to enhance transformer architectures. Through the integration of biologically inspired mechanisms, our model not only matches but also surpasses the efficiency and stability of existing artificial transformer models. By refining synaptic modulation with astrocytic feedback and Hebbian plasticity, we introduce a more adaptive and nuanced attention mechanism that proves advantageous in tasks such as sentiment analysis, image classification, and natural language generation.
While previous models, like the Linearized Transformer inspired from astrocytes [18], also utilize neuron-astrocyte interactions, they lack two critical features: astrocytic non-linearity and relative positional encoding. These omissions limit the model’s ability to capture complex relationships in sequential data, resulting in slower convergence and reduced accuracy and, in some cases, failure to converge entirely—as observed in the Wikitext-2 task. Table II compares the vanilla transformer, linear transformer, linearized transformer inspired by astrocytes, and the proposed Astromorphic Transformer. It maps each model to relevant equations and figures, emphasizing how computational neuroscience equations underpin the bioplausible mechanisms and advancements of our approach.
TABLE II: Comparison of Transformer Architectures
Model Description
Vanilla Transformer Lacks all bioplausibility, employs softmax attention instead of feature maps, and has quadratic complexity for self-attention computation.
Linear Transformer [37] Lacks bioplausibility, employs linear feature maps (replacing softmax), achieving linearized complexity for self-attention (see Fig. S2). Does not incorporate tripartite synapse-based formulation.
Linearized Transformer Inspired from Astrocytes [18] Incorporates bioplausibility and tripartite synapse-based formulation (refer to Eqns. 8, 9, 13, 15, 17 and Fig. 1). Employs linear feature maps (no softmax) and has linearized complexity (Fig. S2). However, it lacks intricate bioplausible plasticities derived from computational neuroscience models.
Astromorphic Transformer (Our Model) Fully bioplausible, employs linear feature maps (no softmax) with linear complexity (Fig. S2) and tripartite synapse formulation (Fig. 1). Shows a clear depiction of the astromorphic self-attention network (Fig. 2) and incorporates astrocytic non-linearity derived from calcium response (refer to Eqns. 5, 14, 16, and Fig. 3). Introduces astrocytic Hebbian plasticities (Eqns. 10, 11) encoding relative positional information (Eqn. 12) inspired by Slow Inward Currents (Eqns. 6, 7). Formulates an accurate model of the Astromorphic Transformer, as detailed in Eqn. 18.
Moreover, our Astromorphic Transformer is catered for simple implementation in current hardware without any additional overhead by leveraging the decomposition of astromorphic attention into distinct write and read modes, making it feasible for traditional computing architectures. Detailed molecular-level computational modeling is not required for practical implementation. Instead, macro-models developed for non-linearity, plasticity, and feedback processes are being utilized, which represent these dynamics in a simplified yet biologically plausible way. These macro-models (see Equation 18) should be easily integrable into conventional hardware, facilitating efficient deployment without needing highly specialized configurations. Energy-efficient in-memory processing with CMOS or emerging post-CMOS devices for transformer architectures [76] can also be potentially utilized for hardware implementation of Astromorphic Transformers. The Astromorphic Transformer can further be improved by incorporating more complex dynamics from neuroscience, such as astrocyte-astrocyte communication, which could enable more sophisticated processing and parallelization of plasticity mechanisms. Additionally, scaling the model to deeper architectures while leveraging characteristics like memory and learning from astrocytes may allow the model to adapt to more complex tasks with long-context sequences.
Acknowledgments
Research was sponsored in part by the Army Research Office and was accomplished under Grant Number W911NF-24-1-0127. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the Army Research Office or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein. This work was also supported partially by the National Science Foundation under award No. EFRI BRAID #2318101.
@@ V-A Appendix A: Transformer Illustration
Transformer employs only encoder architecture for performing sentiment classification and vision tasks. As illustrated in Fig. S1, the input embeddings, combined with the positional embeddings, enter the encoder layer as input tokens. Then key, query, and value are generated from the input tokens, and after performing self-attention on them, a skip connection is added from the inputs, and the overall result is layer normalized and fed into a feed-forward network. Another add & normalize operation is performed before feeding them into a linear layer and softmax function to obtain the output probabilities. Sometimes, there might be more than one encoder layer. For different tasks, like machine translation and text generation, encoder-decoder or only decoder-type architecture of transformers can be used. However, independent of the type of architecture needed, self-attention is a fundamental mechanism defining transformer operation. The “Multi-Head Attention” Block has self-attention modules parallelly running on different heads obtained from the input embeddings.
@@ V-B Appendix B: Memory Efficient Linearized Attention
Linearized attention involves utilizing feature maps with lower complexity levels, thereby minimizing memory and computational resources. As discussed before, the introduction of non-linearity in biological neurons is facilitated by an activation function ($\phi(.)=elu(.)+1$) in the hidden layer neurons. The utilization of this positive similarity feature map as an activation function characterizes the implementation of the linearized attention operation as described by Equation 1. The utilization of linearized feature attention results in a reduction of computational complexity associated with the computation of the attention score. The advantage of utilizing feature map-based linearized attention over the conventional distribution function (i.e. softmax) based self-attention is demonstrated in Fig. S2. Usual self-attention employs softmax as the distribution function; however, the dot product of Query ($Q$) and Key ($K$) poses an $O(N^{2}D)$ complexity, which overuses the memory. Whereas, in linearized attention, the computation multiplies Key ($K$) and Value ($V$), yielding an $O(ND^{2})$ complexity, properly utilizing the memory and computational efficiency when token length ($N$) is larger than internal embedding size ($D$)(We set token embedding dimension equal to the internal embedding size ($d=D$) in this work). Thus, the employment of the linearized feature map enables the realization of memory-efficient linearized attention.
@@ V-C Appendix C: Hyper-parameter Details
Hyper-parameters for the encoder-type astromorphic transformer for IMDB dataset are in Table S1, for the encoder-type convolution and astromorphic self-attention-based transformer for CIFAR10 dataset are in Table S2, and for the astromorphic transformer for Wikitext-2 dataset are in Table S3.
TABLE S1: Hyper-parameter list for the IMDB dataset.
Hyper-parameters Value Tested Range
Non-linearity parameter $\alpha$ $0.25$ $0.15$-$0.4$
Model size $D$ $128$ $32$-$256$
Scaling parameter $m$ $500D$ ($1$-$1000$)$D$
Sequence length $400$ $100$-$600$
Learning rate $0.00001$ $1e^{-4}$ $1e^{-6}$
Batch size $64$ $8$-$256$
Number of epochs $120$ $40$-$150$
Number of heads $4$ $1$-$12$
TABLE S2: Hyper-parameter list for the CIFAR10 dataset.
Hyper-parameters Value Tested Range
Non-linearity parameter $\alpha$ $0.25$ $0.15$-$0.4$
Model size $D$ $256$ $64$-$512$
Scaling parameter $m$ $500D$ ($1$-$1000$)$D$
Patch size $2$ $2$-$4$
Learning rate $0.001$ $1e^{-2}$ $1e^{-4}$
Batch size $32$ $8$-$128$
Number of epochs $250$ $80$-$300$
Number of heads $8$ $2$-$12$
TABLE S3: Hyper-parameter list for the Wikitext-2 dataset.
Hyper-parameters Value Tested Range
Non-linearity parameter $\alpha$ $0.25$ $0.1$-$0.3$
Model size $D$ $768$ $256$-$1024$
Scaling parameter $m$ $100D$ ($50$-$500$)$D$
Sequence length $512$ $128$-$512$
Learning rate $0.00001$ $1e^{-4}$ $5e^{-6}$
Batch size $24$ $8$-$64$
Number of epochs $50$ $20$-$100$
Number of heads $6$ $4$-$12$
V-D Appendix D: Computational Model of Neuron-Astrocyte-Mediated Synaptic Dynamics
In the context of astromorphic computation, astrocytes play a pivotal role in modulating synaptic transmission through both direct and indirect pathways. These pathways, influenced by endocannabinoid signaling and astrocytic calcium dynamics, are essential for the dynamic control of synaptic weights in neuromorphic architectures such as the Astromorphic Transformer.
V-D1 Endocannabinoid Dynamics
Endocannabinoids, such as 2-arachidonoylglycerol (2-AG), are released from the postsynaptic neuron following depolarization. They serve as retrograde messengers, modulating synaptic transmission probability through both depressive and potentiating effects. The release dynamics of 2-AG can be modeled as follows:
$\frac{d(AG)}{dt}=-\frac{AG}{\tau_{AG}}+r_{AG}\cdot\delta(t-t_{\text{spike}})$
where $\tau_{AG}$ is the decay constant of 2-AG, and $r_{AG}$ is the rate of 2-AG production after a spike.
V-D2 Direct Signaling Pathway—Depolarization-Induced Suppression of Excitation (DSE)
In the direct pathway, 2-AG binds to CB1 receptors (CB1Rs) on the presynaptic neuron, reducing the synaptic transmission probability $PR$. This localized effect, termed Depolarization-Induced Suppression of Excitation (DSE), can be described as:
$DSE(t)=-AG(t)\cdot K_{AG}$
where $K_{AG}$ is a constant representing the strength of suppression.
V-D3 Indirect Signaling Pathway—Endocannabinoid-Mediated Synaptic Potentiation (e-SP)
In the indirect pathway, 2-AG binds to CB1Rs on astrocytes, triggering intracellular calcium release from the endoplasmic reticulum (ER) via IP3 receptors (IP3Rs). This calcium release induces astrocytic glutamate release, which potentiates synaptic transmission at distant synapses. The potentiation is given by:
$eSP(t)=m_{eSP}\cdot Glu(t)$
where $m_{eSP}$ is the scaling factor for synaptic potentiation, and $Glu(t)$ is the concentration of astrocytic glutamate.
V-D4 Calcium Dynamics in Astrocytes
The astrocytic calcium response, which regulates synaptic potentiation, is governed by the following equation:
$\frac{d[Ca^{2+}]}{dt}=J_{\text{chan}}(IP3,Ca^{2+})+J_{\text{leak}}-J_{\text{pump}}$
where $J_{\text{chan}}(IP3,Ca^{2+})$ represents calcium influx through IP3-regulated channels (dependent on IP3 and calcium concentration), $J_{\text{leak}}$ accounts for passive calcium leakage from the ER, and $J_{\text{pump}}$ describes calcium reuptake into the ER via ATP-dependent pumps. Once calcium levels exceed a threshold, the astrocyte releases glutamate, governed by:
$\frac{d[Glu]}{dt}=-\frac{Glu}{\tau_{Glu}}+r_{Glu}\cdot H(Ca^{2+}(t)-\theta_{Ca})$
where $\tau_{Glu}$ is the decay constant of glutamate, and $\theta_{Ca}$ is the calcium threshold that triggers release, and $H(x)$ is the Heaviside function defined as $H(x)=1$ if $x\geq 0$ and $H(x)=0$ if $x<0$.
V-D5 Modulation of Synaptic Transmission Probability (PR)
The total synaptic transmission probability $PR(t)$ is influenced by both the direct depressive effect and indirect potentiation, as follows:
$PR(t)=PR(0)+PR(0)\cdot\frac{DSE(t)+eSP(t)}{100}$
where $PR(0)$ is the baseline synaptic transmission probability.
This appendix presents a detailed mathematical framework for neuron-astrocyte-mediated synaptic dynamics, which are key to the functionality of astromorphic computational systems.
V-E Appendix E: Algorithm and Pseudo-code
The pseudo-code (Algorithm 1) represents the core steps of the astromorphic transformer, where neuron-astrocyte interactions modulate synaptic weights, enhancing the self-attention mechanism. The algorithm operates through two phases: Write Mode and Read Mode, followed by the generation of the final transformer output. In the initialization step, the weight matrices for neurons and astrocytes, the presynaptic plasticity parameter, the non-linearity parameter, and the edge matrix representing token positions are set up. During the Write Mode, keys and values are generated from the input tokens using learnable weight matrices. The activation function is applied to introduce non-linearity, resulting in an updated representation. Hebbian learning is then used to update both neuron and astrocyte weight matrices based on the outer product of the activation and the values, while the presynaptic plasticity parameter is updated according to the sum of activated keys modulated by a non-linearity parameter. In the Read Mode, after all tokens have been processed, a query matrix is generated from the input tokens. The activation function is applied to these queries, and the astrocytic calcium response is computed. The modulation factor, which is inversely proportional to the calcium response, adjusts the synaptic weights dynamically. The final output is calculated by matrix multiplication of the synaptic weights and the activated queries along with an addition of a residual connection from the input. The final transformer output is processed through layer normalization to stabilize and accelerate training. A feed-forward network introduces additional non-linearity, followed by another layer normalization step with a residual connection to ensure gradient flow. Finally, a linear transformation and softmax function produce the final logits, representing the output probabilities of the model.
Algorithm 1 Astromorphic Transformer Algorithm
1:  Initialization: Initialize parameters and hyper-parameters
2:  Initialize input tokens $X$, token embedding dimension $d$, hidden layer neurons $m$
3:  Initialize Hebbian weight matrix $H_{neuron}$ and astrocyte weight matrix $H_{astro}$ to 0
4:  Initialize presynaptic plasticity parameter $g$ to 0
5:  Set non-linearity parameter, $\alpha$
6:  Set edge between input tokens $x_{i}$ and $x_{j}$: $a_{ij}$
7:  Write Mode: Encode keys and values from input tokens
8:  for each token $x_{t}$ in input tokens $X$ do
9:     $k_{t}\leftarrow x_{t}W_{k}$
10:     $v_{t}\leftarrow x_{t}W_{v}$
11:     $h_{t}\leftarrow\phi(k_{t})$
12:     {Hebbian learning for neuron and astrocyte plasticity}
13:     $H_{neuron,t}\leftarrow H_{neuron,0}+\frac{1}{m}h_{t}^{T}v_{t}$
14:     $W_{astro}\leftarrow\phi(a_{ij}^{T})$
15:     $H_{astro,t}\leftarrow H_{astro,0}+\frac{1}{m}W_{astro}v_{t}$
16:     {Presynaptic plasticity based on calcium dynamics}
17:     $g\leftarrow g+(\sum(h_{t}))^{\alpha}$
18:  end for
19:  Read Mode: Retrieve information using query for all tokens
20:  $Q\leftarrow XW_{q}$
21:  $h\leftarrow\phi(Q)$
22:  $C\leftarrow hg^{T}$
23:  $P\leftarrow\frac{1}{C}$
24:  $L\leftarrow h((H_{neuron}+H_{astro})\odot P)+X$
25:  $L_{norm}\leftarrow\mathrm{LayerNorm}(L)$
26:  $L_{output}\leftarrow\mathrm{LayerNorm}(\mathrm{FFN}(L_{norm})+L_{norm})$
27:  $logits\leftarrow\mathrm{Softmax}(\mathrm{Linear}(L_{output}))$
28:  return  logits